# Dynamics of Streamers and Pseudostreamers and Implications for the Solar Wind

### v1

@@ Abstract
The origin of the Sun’s slow wind and its inherent variability remain unknown, but there is increasing evidence that interaction between closed and open magnetic flux in the corona plays a role. This paper studies the dynamic evolution of streamers and pseudostreamers with a particular focus on the release of plasma from the closed to the open field region. We employ a global magnetohydrodynamic model that extends from the solar chromosphere to 30 solar radii, and that extends previous interchange magnetic reconnection modelling by including self-consistent thermodynamics. We find that at both the helmet streamer and pseudostreamer there is a continual dynamic interaction between closed and open flux. At the helmet streamer, this takes the form of a “breathing” cycle in which the closed flux contracts and expands, and plasmoids are released along the heliospheric current sheet. The pseudostreamer exhibits a back-and-forth motion, driving interchange reconnection alternately on its opposite flanks. The resulting release of hot, dense plasma leads to density fluctuations in the open field that are significantly larger above the helmet streamer due to the persistence of the plasmoids there. Our model demonstrates that plasma is continually being released into the heliosphere from both streamers and pseudostreamers.
software: Paraview(paraview2011)
@@ 1 Introduction
The solar wind is a stream of plasma that continuously outflows from the Sun and fills the heliosphere. The existence of the solar wind was first proposed by parker1958 with a steady-state, isothermal model. Later, the Mariner 2 spacecraft confirmed the signature of the solar wind using in-situ measurements (neugebauer1962). The observations made from Ulysses satellite further reveal that the solar wind has two different wind streams classified as “fast” and “slow” solar wind (e.g., mccomas2008). The fast solar wind originates mostly from the the central regions of coronal holes where magnetic fields are open (i.e., the magnetic field lines from these regions extend far out into the heliosphere). It has a typical radial speed higher than 500 km s-1 when measured around 1 AU. In recent times, the Solar Orbiter and Parker Solar Probe missions have provided a plethora of information on wind composition and variability using limb observations and in-situ measurements, respectively. For a review relating to the solar wind and its connection to the Sun’s corona, see cranmer2019.
@@
In this paper, we undertake modelling aimed at increasing our understanding of the slow wind. A detailed review of observations and theories relating to the slow wind is given by abbo2016. The precise origin of the streams of slow wind remains to be determined, but there is growing evidence that they typically originate from the peripheries of coronal holes (often around active regions; e.g., owens2013; brooks2015), where open magnetic flux abuts closed magnetic flux. Thus, one proposed mechanism for the formation of the slow wind involves the interaction of closed and open magnetic flux through interchange magnetic reconnection (Suess96; fisk1998; antiochos2011). As shown by antiochos2011, open magnetic flux passing close to closed flux in the low corona can fill a substantial portion of the heliosphere as the so-called “S-Web”. This S-Web is associated with different magnetic structures in the corona that are consistent with the helmet streamers and pseudostreamers found in observations (e.g., scott2018). Recently, wilkins2025 demonstrated that the percentage of heliospheric flux that is rooted in the photosphere within 25 Mm of the closed field varies between around 90% at solar maximum and around 40% at solar minimum, consistent with the solar-cycle variation of fast and slow wind fractions (tokumaru2021).
@@
Computational models of the solar wind are based on a range of different physical models, depending on the target physics. The goal is to provide sufficient thermal pressure in the low corona to drive a solar wind outflow, with some desired characteristics that match observation, such as the radial outflow speed. For all models on these global scales, it is impossible to resolve the (yet-to-be-determined) processes that heat the open and closed magnetic field regions – such as nanoflare reconnection or wave damping mechanisms. Instead, some source term is usually added to the energy equation. One common approach (previously implemented, for example, in the AMRVAC code that we use in this paper) is to control the wind speed by setting a target temperature and adding (or subtracting) energy from the plasma to maintain this temperature. This approach has been used, for example, to study propagation of coronal mass ejections (CMEs), where the goal is simply to have a wind profile (as a function of radius and latitude) that resembles observations to transport these CMEs outwards (e.g., jacobs2005). Other computational approaches seek to reproduce the observed properties of the global corona and solar wind in more detail. For example, lionello2009s present global MHD simulations wherein they seek to reproduce the emission from the Sun across a range of EUV and X-ray wavelengths. They compared a number of different heating terms including a term decaying exponentially with radius to provide a background heating that – when coupled with an additional pressure term in the momentum equation – can provide a fast wind outflow. They also included two magnetic field-dependent terms to account for Quiet Sun and Active Region heating. The latter provides extra heating in strong field regions to match the observed intensity enhancement in active regions (see also downs2010). As well as these empirical heating models, some studies use physically motivated heating models, such as those based on Alfvén wave turbulence (e.g. sokolov2013; vanderholst2014; reville2020). One advantage of such models is that they can reproduce the different heating rates in closed and open field regions self-consistently. Finally, while all of the above-mentioned models have a lower (radial) boundary somewhere above the solar surface, iijima2023 presented a model for an angular wedge that included the sub-surface convection zone out to the solar wind.
@@
Our goal here is to explore the dynamics of helmet streamers and pseudostreamers, particularly in the context of the release of plasma from the closed field to the open field by interchange reconnection. We build on previous work studying the plasma and magnetic field dynamics in the vicinity of the open-closed boundary, which has addressed the dynamics within narrow corridors of open flux (higginson2017b), the motion of isolated parasitic polarity regions (edmondson2010; pontin2013), the formation of a pseudostreamer in the open field by migration from the closed field (scott2021), and the impact of supergranular driving at helmet streamer and pseudostreamer boundaries (aslanyan2021; aslanyan2022). Most of these studies were undertaken using the ARMS code, and all were restricted to an isothermal atmosphere/wind. In particular, aslanyan2021; aslanyan2022 showed that photospheric supergranulation flows drive the formation of interchange-reconnected field lines in narrow filaments. The implication is that the outflow of closed-field plasma should form narrow channels even away from the photosphere, consistent with the imprint of the supergranulation detectable out in the solar wind (borovsky2016). In this paper our goal is to present a model that extends these studies by including a self-consistent thermodynamic evolution of the plasma. This is critical for comparing plasma signatures in the model with those observed by, for example, Solar Orbiter and Parker Solar Probe.
@@
The structure of the paper is as follows. In Section 2 we describe the physics included in our model and the computational method. In Sections 34 we present the results of our simulations. We finish in Section 6 with a discussion.
2 Method
2.1 Governing equations and algorithms
@@
We solve the MHD equations using the MPI-AMRVAC code (porth2014; xia2018; keppens2023). MPI-AMRVAC is a highly modular finite volume Magnetohydrodynamics framework, suited for solving hyperbolic and parabolic partial differential equations associated with various astrophysical phenomena and beyond (keppens2020). In our numerical setup, we define the domain with a 2.5D spherical-polar mesh that allocates all variables in a constant azimuthal plane but includes three components of all vector fields. The radial extent of the domain ranges from the low solar chromosphere, located at $1R_{\odot}$, to the outer solar corona at $30R_{\odot}$. The locations of the grid cell centers are specified on a radially stretched mesh, which increases in radial separation following a geometric progression (xia2018). Further, we utilize seven levels of adaptive mesh refinement on the base stretched grid to include extremely small and large scales at the same time. Our refinement strategy is similar to a Lohner-type prescription (Lohner1987), which is based on computing second-order derivatives weighted by specified variables. For our science case, we set 30% weight on each radial and poloidal component of the magnetic field and 20 % weight on each density and internal energy variable. When mesh refinement occurs, it acts in both the radial and poloidal directions to preserve the primary aspect ratio of the cell. In our domain, the smallest and largest radial grid spaces are approximately $250$ km and $0.78~R_{\odot}$, respectively, under the adaptive mesh refinement setting. The poloidal range linearly covers from the north to the south polar region, with a maximum poloidal grid spacing of $2.81^{\circ}$ and a minimum of $0.02^{\circ}$.
@@
$\frac{\partial e}{\partial t}+{\bf\nabla}.(e\mbox{$u$}{})+p_{g}{\bf\nabla}.\mbox{$u$}{}=H-{\bf\nabla}.\mathbf{Q}-n^{2}\Lambda(T),$ (4)
where $\rho$ is the mass density, $\bf u$} is the plasma velocity, $p_{g}$ is the gas pressure, $p_{m}$ is the magnetic pressure, $\bf B$} is the magnetic field and $e$ is the internal energy density. For the equation of state, we use the Ideal gas law to define $p_{g}$. $\mathbf{g}$ represents the solar gravitational acceleration, which has a maximum value of 274274m s-2 at the lower boundary and follows a $1/r^{2}$ profile with radius. $H$ and $\bf Q$} refer to the coronal heating source and anisotropic conductive heat flux, respectively. Both terms are described comprehensively in Section 2.3. The effect of the optically thin radiative cooling is included by the $n^{2}\lambda(T)$ term, where $n$ represents electron number density and $\lambda(T)$ is the radiative cooling function. We selected a cooling function based on solar coronal abundances (dere2009).
To solve the coupled PDEs in the conservative approach, we choose an HLL-type approximate Riemann solver to discretize the numerical domain and a third-order accurate three-stage Runge-Kutta method for the time integration scheme (Gottlieb1998). The magnetic field is written as the sum of the initial magnetic field plus a (general, non-linear) perturbation, with only the perturbation field being evolved. This approach ensures better numerical stability – particularly in the low plasma-$\beta$ regime (Tanaka1994) – and the associated modifications in governing equations are described thoroughly in xia2018. The spatial reconstruction from cell center to cell interface for the flux computation is performed using the Vanleer limiter (vanleer1974) with second-order accuracy. To maintain the divergence-free constraint for the magnetic field, we opt for both the diffusion (keppens2003) and additional source term approach (powell1999) in our model.
2.2 Initial conditions and boundary conditions
Our aim is to model a system in which the plasma in the magnetically closed regions is hotter and denser than in the open field regions, as observed on the Sun. Because the identities of closed and open field lines will change over time, this should occur self-consistently within the model. This necessitates the presence of a chromospheric layer at the base of our domain, in order to supply the necessary evaporative upflows. Thus, our initial temperature profile starts from 24,500 K at the chromosphere, then includes a transition region and a 1.25 MK hot corona.
@@ 2.2 Initial conditions and boundary conditions
In this model, our goal is to study the helmet streamer and pseudostreamer structures in the solar corona. Thus, to initialise the magnetic field ($\mathbf{B}_{\rm ini}$) in the simulations, we take a superposition of a global dipolar field ($\mathbf{B}_{\rm dip}$) with an additional field ($\mathbf{B_{\rm add}}={\bf\nabla}\times\mathbf{A_{\rm add}}$) only for the following co-latitude($\theta$) range;
$\mathbf{B}_{\rm ini}=\mathbf{B}_{\rm dip}+\mathbf{B_{\rm add}};\quad 58.48^{\circ}\leq\theta\leq 144.43^{\circ}$ (6)
For other $\theta$ values, only the global dipole ($\mathbf{B}_{\rm dip}$) component contributes to the initial field ($\mathbf{B}_{\rm ini}$).
@@
where $B_{0}=-3~G\cdot R^{3}_{\odot}$. The additional magnetic field is defined in terms of the vector potential ($\mathbf{A_{\rm add}}$(talpeanu2022), via
@@
where $A_{0}=2.60~G\cdot R^{5}_{\odot}$, $\theta_{0}=11.45^{\circ}$ and $\delta a=42.97^{\circ}$. At the lower radial boundary, the density and temperature are fixed to their initial values. The bottom boundary is closed for any flows. Both the mass density and radial momentum density are proportional to $r^{-2}$ at the outer radial boundary, where the temperature follows $1/r$ profile. The magnetic field is set as line-tied at the lower radial boundary. At the top boundary, radial, poloidal and azimuthal components of the total magnetic field follow $1/r^{2}$, constant and $1/r$ profiles, respectively.
2.3 Energetics
The thermal heat transport in our model is determined in general by the magnetic field-aligned thermal conduction coefficient ($k_{\parallel}$). For the collisional plasma regime, it is defined in Equ. 10. Where $k_{\parallel}=10^{-6}T^{2.5}~{\rm erg\cdot cm^{-1}\cdot s^{-1}{\cdot K^{-1}}}$. To mimic the effect of reduced collisionality with increasing radius, the conduction coefficients are damped following a $1/r^{2}$ dependency  (shoda2019; matsumoto2021). This is an effectively similar approach to considering the free stream conduction flux for the non-collisonal plasma medium, as the heat is mostly trapped within plasma elements in the absence of strong conduction and the conduction speed is determined by the motion of the plasma elements. Specifically, we let the heat flux($\bf Q$}) in Equation (4) take the form

### v2

@@ Abstract
The origin of the Sun’s slow wind and its inherent variability remain unknown, but there is increasing evidence that interactions between closed and open magnetic flux in the corona play a major role. This paper studies the dynamic evolution of streamers and pseudostreamers with a particular focus on the release of plasma from the closed to the open field region. We employ a global 2.5D magnetohydrodynamic model that extends from the solar chromosphere to 30 solar radii, and that extends previous interchange magnetic reconnection modelling by including self-consistent thermodynamics. We find that at both the helmet streamer and pseudostreamer there is a continual dynamic interaction between closed and open flux even in the absence of explicit driving. At the helmet streamer, the dynamics take the form of a “breathing” cycle in which the closed flux contracts and expands, and plasmoids are released along the heliospheric current sheet. The pseudostreamer exhibits a back-and-forth motion, driving interchange reconnection alternately on its opposite flanks. The resulting release of hot, dense plasma leads to density fluctuations in the open field that are significantly larger above the helmet streamer due to the persistence of the plasmoids there. Our model demonstrates that plasma is continually being released into the heliosphere from both streamers and pseudostreamers. We discuss the implications of our simulation results for observations of the corona and inner heliosphere.
software: MPI-AMRVAC 3.2 (Keppens et al., 2023), Paraview (Numerous, 2011)
@@ I Introduction
The solar wind is a stream of plasma that continuously outflows from the Sun and fills the heliosphere. The existence of the solar wind was first proposed by Parker (1958) with a steady-state, isothermal model. Later, the Mariner 2 spacecraft confirmed the signature of the solar wind using in-situ measurements (Neugebauer and Snyder, 1962). The observations made from satellites such as Ulysses further reveal that the solar wind has two different wind streams classified as “fast” and “slow” solar wind (e.g., McComas et al., 2008). The fast solar wind originates mostly from the the central regions of coronal holes where magnetic fields are open (i.e., the magnetic field lines from these regions extend far out into the heliosphere). It has a typical radial speed greater than 500 km s-1 when measured around 1 AU. In recent times, the Solar Orbiter and Parker Solar Probe (PSP) missions have provided a plethora of information on wind composition and variability using limb observations and in-situ measurements, respectively. For a review relating to the solar wind and its connection to the Sun’s corona, see Cranmer and Winebarger (2019).
@@
In this paper, we undertake modeling aimed at increasing our understanding of the slow wind. A detailed review of observations and theories relating to the slow wind is given by Abbo et al. (2016). The precise origin of the streams of slow wind remains to be determined, but there is growing evidence that they typically originate from the boundaries of coronal holes (often around active regions; e.g., Owens et al., 2014; Brooks et al., 2015; Arge et al., 2024), where open magnetic flux abuts closed magnetic flux. Thus, one proposed mechanism for the formation of the slow wind involves the interaction of closed and open magnetic flux through interchange magnetic reconnection (Suess et al., 1996; Fisk et al., 1998; Antiochos et al., 2011). As shown by Antiochos et al. (2011), open magnetic flux passing close to closed flux in the low corona can fill a substantial portion of the heliosphere as the so-called “S-Web”. This S-Web is associated with different magnetic structures in the corona that are consistent with the helmet streamers and pseudostreamers found in observations (e.g., Scott et al., 2018). Recently, Wilkins et al. (2025) demonstrated that the percentage of heliospheric flux that is rooted in the photosphere within 25 Mm of the closed field varies between around 90% at solar maximum and around 40% at solar minimum, consistent with the solar-cycle variation of fast and slow wind fractions (Tokumaru et al., 2021).
@@
Computational models of the solar wind are based on a range of different physical models, depending on the target physics. The goal is to provide sufficient thermal pressure in the low corona to drive a solar wind outflow, with some desired characteristics that match observations, such as the radial outflow speed. For all models on these global scales, it is impossible to resolve the (yet-to-be-determined) processes that heat the open and closed magnetic field regions – such as nanoflare reconnection or wave damping mechanisms. Instead, some source term is usually added to the energy equation. One common approach (previously implemented, for example, in the MPI-AMRVAC code that we use in this paper) is to control the wind speed by setting a target temperature and adding (or subtracting) energy from the plasma to maintain this temperature. This approach has been used, for example, to study propagation of coronal mass ejections (CMEs), where the goal is simply to have a wind profile (as a function of radius and latitude) that resembles observations in order to transport these CMEs outwards (e.g., Jacobs et al., 2005). Other computational approaches seek to reproduce the observed properties of the global corona and solar wind in more detail. For example, Lionello et al. (2009) present global MHD simulations wherein they seek to reproduce the emission from the Sun across a range of EUV and X-ray wavelengths. They compared a number of different heating terms including a term decaying exponentially with radius to provide a background heating that – when coupled with an additional pressure term in the momentum equation – can provide a fast wind outflow. They also included two magnetic field-dependent terms to account for Quiet Sun and Active Region heating. The latter provides extra heating in strong field regions to match the observed intensity enhancement in active regions (see also Downs et al., 2010). As well as these empirical heating models, some studies use physically motivated heating models, such as those based on Alfvén wave turbulence (e.g. Sokolov et al., 2013; van der Holst et al., 2014; Réville et al., 2020). One advantage of such models is that they can reproduce the different heating rates in closed and open field regions self-consistently. Finally, while all of the above-mentioned models have a lower (radial) boundary somewhere above the solar surface, Iijima et al. (2023) presented a model for an angular wedge that included the sub-surface convection zone out to the solar wind.
@@
Our goal here is to explore the dynamics of helmet streamers and pseudostreamers, particularly in the context of the release of plasma from the closed field to the open field by interchange reconnection. We build on previous work studying the plasma and magnetic field dynamics in the vicinity of the open-closed boundary, which has addressed the dynamics within narrow corridors of open flux (Higginson et al., 2017), the motion of isolated parasitic polarity regions (Edmondson et al., 2010; Pontin et al., 2013), the formation of a pseudostreamer in the open field by migration from the closed field (Scott et al., 2021), and the impact of supergranular driving at helmet streamer and pseudostreamer boundaries (Aslanyan et al., 2021, 2022). Most of these studies were undertaken using the ARMS code, and all were restricted to an isothermal atmosphere/wind. In particular, Aslanyan et al. (2021, 2022) showed that photospheric supergranulation flows drive the formation of interchange-reconnected field lines in narrow filaments. The implication is that the outflow of closed-field plasma should form narrow channels even away from the photosphere, consistent with the imprint of the supergranulation detectable out in the solar wind (Borovsky, 2016). In this paper our goal is to present a model that extends these studies by including a self-consistent thermodynamic evolution of the plasma. This is critical for comparing plasma signatures in the model with those observed by, for example, Solar Orbiter and Parker Solar Probe.
@@
The structure of the paper is as follows. In Section II we describe the physics included in our model and the computational method. In Sections IIIIV we present the results of our simulations. In Section V we compute synthetic observable and discuss the relevance for observations, and in Section VI we finish with a discussion.
II Method
II.1 Governing equations and algorithms
@@
We solve the MHD equations using the MPI-AMRVAC code (Porth et al., 2014; Xia et al., 2018; Keppens et al., 2023). MPI-AMRVAC is a highly modular finite volume Magnetohydrodynamics framework, suited for solving hyperbolic and parabolic partial differential equations associated with various astrophysical phenomena and beyond (Keppens et al., 2020). In our numerical setup, we define the domain with a 2.5D spherical-polar mesh that allocates all variables on a constant azimuthal plane but includes three components of all vector fields. The radial extent of the domain ranges from the low solar chromosphere, located at $1R_{\odot}$, to the outer solar corona at $30R_{\odot}$. The locations of the grid cell centers are specified on a radially stretched mesh, which increases in radial separation following a geometric progression (Xia et al., 2018). Further, we utilize seven levels of adaptive mesh refinement on the base stretched grid to include extremely small and large scales at the same time. Our refinement strategy is similar to a Lohner-type prescription (Lohner, 1987), which is based on computing second-order derivatives weighted by specified variables. Specifically, we set 30% weight on each radial and poloidal component of the magnetic field and 20 % weight each on the density and internal energy. When mesh refinement occurs, it acts in both the radial and poloidal directions to preserve the primary aspect ratio of the cell. In our domain, the smallest and largest radial grid spaces are approximately $250$ km and $0.78~R_{\odot}$, respectively, under the adaptive mesh refinement setting. The poloidal range linearly covers from the north to the south polar region, with a maximum poloidal grid spacing of $2.81^{\circ}$ and a minimum of $0.02^{\circ}$.
@@
$\frac{\partial e}{\partial t}+{\bf\nabla}\cdot(e\mbox{$u$}{})+p_{g}{\bf\nabla}\cdot\mbox{$u$}{}=H-{\bf\nabla}\cdot\mathbf{Q}-n^{2}\Lambda(T),$ (4)
where $\rho$ is the mass density, $\bf u$} is the plasma velocity, $p_{g}$ is the gas pressure, $p_{m}$ is the magnetic pressure, $\bf B$} is the magnetic field and $e$ is the internal energy density. For the equation of state, we use the Ideal gas law to define $p_{g}$. $\mathbf{g}$ represents the solar gravitational acceleration, which has a maximum value of 274274m s-2 at the lower boundary and follows a $1/r^{2}$ profile with radius. $H$ and $\bf Q$} refer to the coronal heating source and anisotropic conductive heat flux, respectively. Both terms are described comprehensively in Section II.3. The effect of the optically thin radiative cooling is included by the $n^{2}\lambda(T)$ term, where $n$ represents electron number density and $\lambda(T)$ is the radiative cooling function. We selected a cooling function based on solar coronal abundances (Dere et al., 2009).
To solve the coupled PDEs in the conservative approach, we choose an HLL-type approximate Riemann solver to discretize the numerical domain and a third-order-accurate three-stage Runge-Kutta method for the time integration scheme (Gottlieb and Shu, 1998). The magnetic field is written as the sum of the initial magnetic field plus a (general, non-linear) perturbation, with only the perturbation field being evolved. This approach ensures better numerical stability – particularly in the low plasma-$\beta$ regime (Tanaka, 1994) – and the associated modifications in governing equations are described thoroughly in Xia et al. (2018). The spatial reconstruction from cell center to cell interface for the flux computation is performed using the Vanleer limiter (van Leer, 1974) with second-order accuracy. To maintain the divergence-free constraint for the magnetic field, we opt for both the diffusion (Keppens et al., 2003) and additional source term approach (Powell et al., 1999) in our model.
II.2 Initial conditions and boundary conditions
Our aim is to model a system in which the plasma in the magnetically closed regions is hotter and denser than in the open field regions, as observed on the Sun. Because the identities of closed and open field lines will change over time, this should occur self-consistently within the model. This necessitates the presence of a chromospheric layer at the base of our domain, in order to supply the necessary evaporative upflows. Thus, our initial temperature profile starts from 24,500 K at the chromosphere, then includes a transition region and a 1.25 MK hot corona. At $t=0$ we impose the temperature profile
@@ II.2 Initial conditions and boundary conditions
In this model, our goal is to study helmet streamer and pseudostreamer structures in the solar corona. Thus, to initialise the magnetic field ($\mathbf{B}_{\rm ini}$) in the simulations, we take a superposition of a global dipolar field ($\mathbf{B}_{\rm dip}$) with an additional field ($\mathbf{B_{\rm add}}={\bf\nabla}\times\mathbf{A_{\rm add}}$) that contributes only in the following co-latitude($\theta$) range;
$\mathbf{B}_{\rm ini}=\mathbf{B}_{\rm dip}+\mathbf{B_{\rm add}};\quad 1.02~(58.48^{\circ})\leq\theta\leq 2.52~(144.43^{\circ})$ (6)
Note that $\mathbf{B_{\rm add}}$ has been constructed so that it vanishes at the end points of the co-latitude range above. Outside this $\theta$ range, we take only the global dipole ($\mathbf{B}_{\rm dip}$) component to contribute to the initial field ($\mathbf{B}_{\rm ini}$).
@@
where $B_{0}=-3~G\cdot R^{3}_{\odot}$. The additional magnetic field is defined in terms of the vector potential ($\mathbf{A_{\rm add}}$(Talpeanu et al., 2022), via
@@
where $A_{0}=2.60~G\cdot R^{5}_{\odot}$, $\theta_{0}=0.2~(11.45^{\circ})$ and
$\delta a=0.75~(42.97^{\circ})$. At the lower radial boundary, the density and temperature are fixed to their initial values for the duration of the simulation. The bottom boundary is closed for any flows. Both the mass density and radial momentum density are proportional to $1/r^{2}$ at the outer radial boundary, where the temperature follows a $1/r$ profile. The magnetic field is set as line-tied at the lower radial boundary. At the top boundary, radial, poloidal and azimuthal components of the total magnetic field follow $1/r^{2}$, constant and $1/r$ profiles, respectively.
II.3 Energetics
The thermal heat transport in our model is determined in general by the magnetic field-aligned thermal conduction coefficient ($k_{\parallel}$). For the collisional plasma regime, it is defined in Equ. 10, where $k_{\parallel}=10^{-6}T^{2.5}~{\rm erg\cdot cm^{-1}\cdot s^{-1}{\cdot K^{-1}}}$. To mimic the effect of reduced collisionality with increasing radius, the conduction coefficients are damped following a $1/r^{2}$ dependency  (Shoda et al., 2019; Matsumoto, 2021). This is an effectively similar approach to considering the free stream conduction flux for the non-collisonal plasma medium, as the heat is mostly trapped within plasma elements in the absence of strong conduction and the conduction speed is determined by the motion of the plasma elements. Specifically, we let the heat flux($\bf Q$}) in Equation (4) take the form